SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Trigger Operation Meeting (9/11/07)

Attendees: Rainer Bartoldus, Kim Hojeong, Su Dong, Selina Li, Mark Tibbetts, Al Eisner.

Meeting Information

 Date      :  Tuesday, September 11th, 2007
 Time      :  2:00pm - 3:30pm PST
 SLAC Room :  Madrone conf room, ROB 
 
 Dial-in   :  1-510-665-5437, pass code 6649#

--- Minutes ---

Operations

  • TRG crates are still powered on. It will be powered off when IFR people are done with cosmics runs.
  • Task list review. I left over what we agreed that it's still not too late to do and will be beneficial. Need more discussion on who will do it and when.

EMT

  • EMT uses UPC serial number to compare with expected value. So when UPC is changed, it will cause error which is made by this script, but it won't be noticed by any place else (EMT calibration never fails). Solution: Mark will change script to fail EMT calibration if there are non zero entries in the plots, so that the shifters will call him and he'll take a look.

DCT

GLT

Online

L3

  • Su Dong and Selina's L3EMT filter study (directory).
    new filter has the following cuts:
    1. At least 3 L3 EMC Clusters
    2. No L3 track what so ever
    3. effective mass cut at 800 MeV/c2
    4. reject if biggest energy cluster's energy is bigger than 3.5 GeV (plot)
    Summary plots for data and MC. These plots (data and MC) have all cuts except event mass cut - mostly backgrounds. The reco events are here (data and MC). They used pi0pi0 events below 2 GeV for MC.
  • EMC people noticed a blob problem after 1E implementation. Rainer prepared plots before and after the 1E implementation. The close up plots are: before and after. The high energy of the blob is understandable by 1E. The early timing could be related with the priority of the line (8th): maybe events later in the priority can go through 1E line and shift timing.
  • Arik's code validation result. This validation is not to see if Arik's code makes improvements, but to see if the change without new seed track configuration is still suitable for simulating data taken with the current L3, and how much it will affect monitoring and calibration lines. The changes in code may reduce number of ghost tracks, but it should not change real track numbers. Although the changes in numbers (in result) are in general small, we loose real tracks, in B0 generic MC, L3THadronicAFilter for example. So it's not clear if this is OK.

DQM

  • NTR by Rahmat
  • We need some conclusion on data quality of Run6.

Trickle injection monitoring

Trigger Configuration

  • RecoBGFilter-tagBits (plot) made from real run with 1E in: Yury's invisible mode is a little over 1 nb, using 3~4% rate. He needs to work on it to make less effect. The calibration lines still need work.

AOB

  • Mark will be away from Monday until October 11th.
  • We are going to bi-weekly mode. But the next meeting in two weeks is canceled due to CM. We'll see on October 9th. Please give a talk on what you did during CM parellel session.

This page was created by Kim Hojeong (hojeong@slac); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.

Last Modified: September 12, 2007.